# Emancipatory Information Retrieval

### v4

@@ 5 Conclusion
The field of IR has undergone seismic shifts over the last decade largely instigated by rapid developments in deep learning [94] and more recently generative AI [137]. Between trying to stay updated in the face of daily firehoses of new AI publications and chasing the coveted badges of “state-of-the-art” it can be easy as an IR researcher to overlook the pressing needs of society and envision only the possibilities as conjured by the latest AI news cycle. If we are to address the multitude of crises facing our world today—spanning the ecological, the social, the political, and the economic spectrums—IR must undergo another seismic shift over the next decade. This piece is a provocation for the IR community to situate itself in the generational emancipatory struggles and stand in resistance to the growing epidemic of authoritarianism and fascism through the practices of making and unmaking of information access technologies. It is also an invitation to scholars from other fields as well as to activists and artists to join us and be part of this movement. While focused on IR, it is also a challenge to the broader computing and technology community to reimagine our work to not serve the privileged few by further concentrating immense power and wealth, but as a liberating force that connects us and humanizes us universally. Technology after all is a project of world-making. So, what kind of a world are we collectively going to build together?

### v6

@@ 5 Conclusion
The field of IR has undergone seismic shifts over the last decade largely instigated by rapid developments in deep learning (mitra2018introduction) and more recently generative AI (white2025information). Between trying to stay updated in the face of daily firehoses of new AI publications and chasing the coveted badges of “state-of-the-art” it can be easy as an IR researcher to overlook the pressing needs of society and envision only the possibilities as conjured by the latest AI news cycle. If we are to address the multitude of crises facing our world today—spanning the ecological, the social, the political, and the economic spectrums—IR must undergo another seismic shift over the next decade. This piece is a provocation for the IR community to situate itself in the generational emancipatory struggles and stand in resistance to the growing epidemic of authoritarianism and fascism through the practices of making and unmaking of information access technologies. It is also an invitation to scholars from other fields as well as to social movement organizers and artists to join us and be part of this shared struggle. And as IR researchers, we must remember that “it is vital that we approach these spaces with curiosity and humility; in recognition of our own incomplete understanding of the world; open to change and be changed by these encounters” (mitra2025search).
A central thesis underlying this work is that the emancipatory struggles in the real world and one that we are encouraging the IR community to acknowledge and center within our research are one and the same. So, the work of emancipatory IR and the work of movement building for social justice and emancipation outside of IR must co-construct the conditions for dismantling oppressive regimes, such as those of colonialism, cisheteropatriarchy, and capitalism. While focused on IR, it is also a challenge to the broader computing and technology community to reimagine our work to not serve the privileged few by further concentrating immense power and wealth, but as a liberatory force that connects us and humanizes us universally. Technology after all is a project of world-making. So, what kind of a world are we collectively going to build together?
Positionality statement
I, the author of this paper, spent most of my career at a large technology corporation in the global north. However, the perspectives presented in this work is intended to challenge Big Tech and global north’s view of technology and our collective futures. Outside of information retrieval research, I participate in social movement spaces and Southasian anti-fascist organizing. I stand in solidarity with the peoples of Palestine, Sudan, Congo, Rohingya, and other groups who are being subjected to unimaginable oppression and ongoing genocide. I also stand in solidarity with the people of Kashmir in their emancipatory struggles against state violence. These ongoing atrocities and cases of systemic oppression significantly informed and influenced my work.
Acknowledgments and Disclosure of Funding
The author gratefully acknowledges feedback from Ida Larsen-Ledet on the various drafts of this paper. No external funding was received in support of this work.